# Parameter-Efficient Fine-Tuning via Circular Convolution

### v2

@@ 3.3 Backpropagation
To implement backpropagation efficiently, we require the analytical derivatives of the loss function $\mathcal{L}$, with respect to $\Delta\mathbf{w}$ and $\mathbf{x}$. Utilizing the chain rule, these derivatives can be expressed as:
@@
Since $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$. For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\mathcal{C}(\Delta\mathbf{w})\mathbf{x}=\mathcal{C}(\mathbf{x})\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$. Substituting these results into Equation 2, we obtain:
@@
By explicitly implementing this derivative calculation using Equation 1, backpropagation can also enjoy the computational efficiency provided by the FFT algorithm.
@@ 3.4 Block-Circular Convolution
Despite the elegance and efficiency of the circular convolution operator, it has 2 fundamental limitations due to the requirement that the convolution kernel must be the same size as the activation vector: ❶ It cannot be applied to non-square weight matrices.The number of learnable parameters cannot be altered. The first limitation restricts its application in scenarios such as fine-tuning a LLaMA3-8B model, where the weight matrix sizes include $4096\times 1024$. The second limitation reduces the flexibility of C3A, making it challenging to handle complex downstream tasks that require a larger number of learnable parameters. To address these limitations, we resort to block-circular convolution. By splitting the activation vector $\mathbf{x}$ and the post-transformation vector $\mathbf{z}$ into blocks of the same size, we can assign unique convolution kernels to each pair of these blocks. To be concrete,
$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{\frac{d_{2}}{b}}\end{bmatrix},\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{\frac{d_{1}}{b}}\end{bmatrix},$
@@
$\Delta\mathbf{z}=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{x},\qquad\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1\frac{d_{2}}{b}})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2\frac{d_{2}}{b}})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}1})&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}\frac{d_{2}}{b}})\end{bmatrix}.$
In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{b}$ and we have $\frac{d_{1}d_{2}}{b^{2}}b=\frac{d_{1}d_{2}}{b}$ learnable parameters. In fact, $b$ acts as a hyperparameter controlling the number of learnable parameters, just like $r$ for LoRA. However, we highlight that while $r$ controls the rank of the delta matrix and learnable parameter number at the same time, $b$ only affects the latter. Disentangling these attributes enables higher adaptation flexibility and potential better results.
@@ 3.5.1 Time Complexity
LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(r(d_{1}+d_{2}))$. Typically, $r\ll\min(d_{1},d_{2})$. Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}(d_{1}+d_{2}))$, where $r_{v}$ can exceed $\max(d_{1},d_{2})$. Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
Thanks to the $\mathcal{O}(n\log n)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$. The first term is the time complexity for FFT, and the second term is for aggregation. In practical scenarios, $b$ is chosen as the greatest common divisor of $d_{1}$ and $d_{2}$ to achieve a high compression ratio. Given that, C3A is comparable to LoRA in terms of time complexity.
@@
Method Time # Param # Other # Total
LoRA $\mathcal{O}(r(d_{1}+d_{2}))$ $r(d_{1}+d_{2})$ 0 $r(d_{1}+d_{2})$
VeRA $\mathcal{O}(r_{v}(d_{1}+d_{2}))$ $r_{v}+d_{1}$ $r_{v}(d_{1}+d_{2})$ $r_{v}(d_{1}+d_{2})+r_{v}+d_{1}$
C3A $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$ $\frac{d_{1}d_{2}}{b}$ $b$ $\frac{d_{1}d_{2}}{b}+b$
@@ 3.5.2 Space Complexity
We analyze the space complexity of LoRA, VeRA, and C3A during training. The differences among these methods primarily arise from the trainable parameters and the auxiliary tensors required for the forward pass and backpropagation. LoRA does not rely on auxiliary tensors, while VeRA necessitates 2 random projection matrices, with a total size of $r_{v}(d_{1}+d_{2})$. Since $r_{v}$ is by no means negligible, the memory consumption of VeRA would be significantly larger than that of LoRA.
In terms of C3A, the only additional auxiliary tensor would be of size $b\leq\min(d_{1},d_{2})$, which is reserved by the FFT algorithm. Given a proper $b$, which is often close to the greatest common divisor of $d_{1}$ and $d_{2}$, the space complexity of C3A excels. Additionally, we observe that when applying FFT, input blocks and output blocks are independent within their own group. Therefore, parallelization can be employed to enable the flexibility of trading space for time.
@@ 4 Experiment
In this section, we firstly conduct experiments on a synthetic dataset to show the superior expressiveness of C3A over LoRA. After that, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV). For NLP, we first demonstrate the effectiveness of C3A using the GLUE benchmark with RoBERTa-Base and RoBERTa-Large models. Additionally, we apply C3A to fine-tuning the LLaMA family models, which stands as a representative of prevalent large language models (LLMs). For CV, we examine the most representative classification task using Vision Transformers (ViTs) on a wide range of datasets.
4.1 Synthetic Data
Settings. We distribute 8 points evenly on a 2D plane as cluster centers and randomly sample 30 points from 8 corresponding multivariate Gaussian distributions. A 3-layer MLP is then used to classify these point clusters. To compare the expressiveness of 2 types of layers, we replace the middle layer with either a low-rank layer or a circulant layer, ensuring that both layers have the same number of trainable parameters for a fair comparison.
Figure 3: Expressiveness test on synthetic data. The left figure shows the synthetic data used for the experiment, while the right figure illustrates the training accuracy curves of a 3-layer MLP, incorporating C3A, LoRA, and standard linear layers, respectively.
Results. The results are presented in Figure 3. We observe that LoRAr=1 struggles with this straightforward classification task. In contrast, C3Ab=128/2, despite using the same number of parameters, achieves perfect classification, comparable to a standard linear layer. This demonstrates the high expressiveness of C3A given the same parameter budget.
4.2 Natural Language Understanding
Baselines. We compare our C3A with several representative PEFT methods, including BitFit Zaken et al. (2021), (IA)3 Liu et al. (2022), LoRA Hu et al. (2021), VeRA Kopiczko et al. (2023), and BOFT Liu et al. (2023). BitFit selectively fine-tunes existing parameters, specifically the biases. (IA)3 is the state-of-the-art method that adds extra adapters. LoRA is a widely known PEFT method that employs low-rank decomposition to compress additive delta matrices. VeRA is a recent approach that focuses on further reducing trainable parameters of LoRA while preserving a high rank. BOFT is another innovative method in PEFT research, compressing multiplicative delta matrices using orthogonal decomposition and butterfly factorization.
Table 2: Performance of different PEFT methods on the GLUE benchmark. We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets. We report the Matthew’s Correlation Coefficient (MCC) for CoLA, Pearson Correlation Coefficient (PCC) for STS-B, and accuracy (Acc.) for all the remaining tasks. For each metric, a higher score indicates better performance. “Avg.” denotes the average score of each method across all datasets. The best results for each dataset are highlighted in bold. # Trainable parameters does not include the classification head since each method uses a head of the same size. Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
Method # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
Base Full 124M 17.19 94.01$\pm$0.39 87.10$\pm$0.79 62.00$\pm$1.16 92.40$\pm$0.28 77.33$\pm$2.68 90.70$\pm$0.14 83.92
BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
(IA)3  [^1] K 19.86 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
LoRAr=8 295K 13.75 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
VeRAr=1024 43K 15.51 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
@@ 4.2 Natural Language Understanding
C3Ab=768/1 18K 12.83 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
C3Ab=768/6  [^1] K 12.72 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
Large BitFit 271K 30.65 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
(IA)3 295K 48.81 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
LoRAr=8 786K 34.12 95.53$\pm$0.35 86.12$\pm$3.22 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
VeRAr=256 61K 34.16 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
@@
C3Ab=1024/1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
C3Ab=1024/8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
Settings. We evaluate our proposed C3A on the General Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018), which encompasses a wide range of natural language understanding (NLU) tasks, including single-sentence classification, similarity and paraphrase, and natural language inference. To enhance practicality, we split these datasets following the train-validation-test approach. The best-performing model is selected based on validation set performance across the fine-tuning epochs, and the reported performance corresponds to its performance on the test set. For this evaluation, we fine-tune the pre-trained RoBERTa-Base and RoBERTa-Large models Liu et al. (2019). For the unique hyperparameters of different baselines, we adopt the values suggested in the original papers (e.g., VeRA’s $r$ and BOFT’s $b$ and $m$). The number of trainable parameters excludes the classification head, as each method uses one of the same size. The shared hyperparameters (i.e., the learning rate for classification head and for other trainable parameters, separately) are found by hyperparameter search. For the memory cost, to ensure fairness and consistency, we fix the length of input data to 256 tokens and use a batchsize of 64.

### v3

@@ 3.3 Backpropagation
To effectuate backpropagation with optimal efficiency, it is imperative to obtain the analytical derivatives of the loss function $\mathcal{L}$ with respect to $\Delta\mathbf{w}$ and $\mathbf{x}$. Following the approach outlined in Ding et al. (2017), we aim to explain backpropagation using simpler language. By applying the chain rule, these derivatives are delineated as follows:
@@
Given that $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it logically follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$. Concerning $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we observe the commutative property of the circular convolution operation (i.e., $\mathcal{C}(\Delta\mathbf{w})\mathbf{x}=\mathcal{C}(\mathbf{x})\Delta\mathbf{w}$), which implies $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$. Substituting these findings into Equation 2, we derive:
@@
By meticulously executing this derivative computation in accordance with Equation 1, backpropagation can harness the computational efficacy facilitated by the FFT algorithm.
@@ 3.4 Block-Circular Convolution
Notwithstanding the elegance and efficiency of the circular convolution operator, it is subject to two fundamental limitations stemming from the constraint that the convolution kernel must match the dimensions of the activation vector: ① It is inapplicable to non-square weight matrices.The count of learnable parameters remains fixed. The first restriction hampers its applicability in scenarios such as fine-tuning a LLaMA3-8B model, where the weight matrix dimensions include $4096\times 1024$. The second constraint diminishes the adaptability of C3A, presenting challenges in addressing complex downstream tasks that necessitate a greater number of learnable parameters. To mitigate these limitations, we employ block-circular convolution (Ding et al., 2017). By partitioning the activation vector $\mathbf{x}$ and the post-transformation vector $\Delta\mathbf{z}$ into blocks of identical size, unique convolution kernels can be allocated to each pair of these blocks. Specifically,
$\mathbf{x}$ $=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{\frac{d_{2}}{b}}\end{bmatrix}$
$\Delta\mathbf{z}$ $=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{\frac{d_{1}}{b}}\end{bmatrix},$
@@
$\Delta\mathbf{z}$ $=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{x}$ (3)
$\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})$ $=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1\frac{d_{2}}{b}})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2\frac{d_{2}}{b}})\\ \cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}1})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}\frac{d_{2}}{b}})\end{bmatrix}.$ (4)
We refer our readers to Algorithm A1 in Appendix A for a Pytorch implementation. In this context, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{b}$, and it follows that $\frac{d_{1}d_{2}}{b^{2}}b=\frac{d_{1}d_{2}}{b}$ represents the number of learnable parameters. Notably, the parameter $b$ serves as a hyperparameter modulating the quantity of learnable parameters, analogous to the role of $r$ in LoRA. It is imperative to distinguish, however, that whereas $r$ simultaneously governs the rank of the delta matrix and the number of learnable parameters, $b$ exclusively influences the latter. This disentanglement of matrix rank and parameter count facilitates greater adaptability and potentially yields superior outcomes.
@@ 3.5.1 Time Complexity
LoRA integrates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are successively multiplied with the activation vector, resulting in a computational complexity of $\mathcal{O}(r(d_{1}+d_{2}))$. Generally, $r\ll\min(d_{1},d_{2})$. In contrast, VeRA, despite its high-rank structure and relatively few trainable parameters, suffers from a prohibitive computational complexity of $\mathcal{O}(r_{v}(d_{1}+d_{2}))$, where $r_{v}$ can exceed $\max(d_{1},d_{2})$. Consequently, striking an optimal balance between high rank and computational efficiency remains an elusive task.
On GPUs, the cuFFT backend automatically parallelizes FFT operations along the axes not being transformed, with the degree of parallelism $p$ determined by the available resources. Thanks to the $\mathcal{O}(n\log n)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}(\frac{(d_{1}+d_{2})}{p}\log{b}+\frac{d_{1}d_{2}}{b})$. The first term is the time complexity for FFT and the second term is for aggregation. In practical scenarios, $b$ is chosen as the greatest common divisor of $d_{1}$ and $d_{2}$ to achieve a high compression ratio. Given that, C3A is comparable to LoRA in time complexity.
Method Time Space
# Param # Other
LoRA $\mathcal{O}(r(d_{1}+d_{2}))$ $r(d_{1}+d_{2})$ 0
VeRA $\mathcal{O}(r_{v}(d_{1}+d_{2}))$ $r_{v}+d_{1}$ $r_{v}(d_{1}+d_{2})$
C3A $\mathcal{O}(\frac{d_{1}+d_{2}}{p}\log{b}+\frac{d_{1}d_{2}}{b})$ $\frac{d_{1}d_{2}}{b}$ $pb$
@@ 3.5.2 Space Complexity
We analyze the space complexity of LoRA, VeRA, and C3A during training. The differences among these methods primarily arise from the trainable parameters and the auxiliary tensors required for the forward pass and backpropagation. LoRA does not rely on auxiliary tensors, while VeRA necessitates 2 random projection matrices, with a total size of $r_{v}(d_{1}+d_{2})$. Since $r_{v}$ is by no means negligible, the memory usage of VeRA is significantly larger than that of LoRA.
In terms of C3A, the only additional auxiliary tensor would be of size $pb\leq\min(d_{1},d_{2})$, which is reserved by the FFT algorithm. By selecting an appropriate $b$, which is often close to the greatest common divisor of $d_{1}$ and $d_{2}$, the space complexity of C3A is optimized. Furthermore, because $p$ scales with the available resources, the algorithm inherently manages dynamic memory consumption without additional effort.
@@ 4 Experiment
Methods # Params Mem SST-2 MRPC CoLA QNLI RTE STS-B Avg.
Base Full 124M 17.19G 94.01$\pm$0.39 87.10$\pm$0.79 62.00$\pm$1.16 92.40$\pm$0.28 77.33$\pm$2.68 90.70$\pm$0.14 83.92
BitFit 0.102M 12.60G 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
(IA)3 0.111M 19.86G 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
LoRAr=8 0.295M 13.75G 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
VeRAr=1024 0.043M 15.51G 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
@@
C3Ab=768/1 0.018M 12.83G 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
C3Ab=768/6 0.111M 12.72G 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
Large Full 354M 43.40G 95.75$\pm$0.45 88.35$\pm$0.64 64.87$\pm$1.25 92.40$\pm$0.28 84.48$\pm$1.14 91.65$\pm$0.14 86.25
BitFit 0.271M 30.65G 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
(IA)3 0.295M 48.81G 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
LoRAr=8 0.786M 34.12G 95.53$\pm$0.35 86.12$\pm$0.86 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
VeRAr=256 0.061M 34.16G 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
@@
C3Ab=1024/1 0.049M 31.83G 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
C3Ab=1024/8 0.393M 31.79G 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
Table 2: Performance of different PEFT methods on the GLUE benchmark. We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets. We report the Matthew’s Correlation Coefficient (MCC) for CoLA, Pearson Correlation Coefficient (PCC) for STS-B, and accuracy (Acc.) for all the remaining tasks. For each metric, a higher score indicates better performance. “Avg.” denotes the average score of each method across all datasets. The best results for each dataset are highlighted in bold. # Params does not include the classification head since each method uses a head of the same size. Memory cost (Mem) is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
Baselines.
We compare our C3A with several representative PEFT methods, including BitFit (Zaken et al., 2021), (IA)3 (Liu et al., 2022), LoRA (Hu et al., 2021), VeRA (Kopiczko et al., 2023), BOFT (Liu et al., 2023) and DoRA (Liu et al., 2024b). BitFit fine-tunes biases. (IA)3 is a SOTA method adding adapters. LoRA uses low-rank decomposition to compress additive delta matrices. VeRA reduces LoRA’s trainable parameters while maintaining a high rank. BOFT compresses multiplicative delta matrices via orthogonal decomposition and butterfly factorization. DoRA separately learns magnitude and direction of delta matrices.
Methods Params (%) Mem BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
ChatGPT - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA2-7B LoRAr=32 0.83 41.71G 71.0 81.4 79.6 87.4 83.2 82.6 67.5 81.5 79.3
VeRAr=16384 0.05 57.88G 68.9 $\downarrow 2.1$ 81.0 $\downarrow 0.4$ 77.4 $\downarrow 2.2$ 87.5 $\uparrow 0.1$ 81.3 $\downarrow 1.9$ 81.1 $\downarrow 1.5$ 62.8 $\downarrow 4.7$ 79.9 $\downarrow 1.6$ 77.5 $\downarrow 1.8$
[…]
DoRAr=32 0.84 55.24G 72.3 $\uparrow 1.3$ 84.2 $\uparrow 2.8$ 78.7 $\downarrow 0.9$ 88.5 $\uparrow 1.1$ 84.0 $\uparrow 0.8$ 80.7 $\downarrow 1.9$ 69.1 $\uparrow 1.6$ 83.7 $\uparrow 2.2$ 80.2 $\uparrow 0.9$
C3Ab=4096/32 0.35 44.89G 73.4 $\uparrow 2.4$ 83.4 $\uparrow 2.0$ 82.1 $\uparrow 2.5$ 88.3 $\uparrow 0.9$ 84.9 $\uparrow 1.7$ 86.6 $\uparrow 4.0$ 66.7 $\downarrow 0.8$ 82.5 $\uparrow 1.0$ 81.0 $\uparrow 1.7$
LLaMA3-8B LoRAr=32 0.70 51.18G 73.8 88.2 80.4 94.0 85.5 87.5 78.1 84.0 83.9
VeRAr=16384 0.04 66.03G 72.2 $\downarrow 1.6$ 87.5 $\downarrow 0.7$ 80.0 $\downarrow 0.4$ 94.3 $\uparrow 0.3$ 83.9 $\downarrow 1.6$ 84.5 $\downarrow 3.0$ 75.6 $\downarrow 2.5$ 82.4 $\downarrow 1.6$ 82.5 $\downarrow 1.4$
[…]
DoRAr=32 0.71 63.37G 75.3 $\uparrow 1.5$ 88.9 $\uparrow 0.7$ 82.2 $\uparrow 1.8$ 96.7 $\uparrow 2.7$ 86.0 $\uparrow 0.5$ 89.0 $\uparrow 1.5$ 79.8 $\uparrow 1.7$ 84.9 $\uparrow 0.9$ 85.3 $\uparrow 1.4$
C3Ab=4096/32 0.26 56.08G 76.9 $\uparrow 3.1$ 91.4 $\uparrow 3.2$ 82.1 $\uparrow 1.7$ 94.9 $\uparrow 0.9$ 86.9 $\uparrow 1.4$ 89.6 $\uparrow 2.1$ 79.4 $\uparrow 1.3$ 86.1 $\uparrow 2.1$ 85.9 $\uparrow 2.0$
Table 3: Comparison of various methods on the LLaMA2-7B and LLaMA3-8B models across eight commonsense reasoning datasets. For each language model, the best result on each dataset is shown in bold. The symbols $\uparrow$ and $\downarrow$ indicate relative improvement and decrease, respectively, compared to the LoRA method. Experiments are conducted on a single H800 GPU with 80GB HBM.
Methods Params (%) Math Avg. Code Avg.
GSM8K MATH HumanEval HumanEval+ MBPP MBPP+
LLaMA2-7B Zero-Shot - 7.3 1.1 4.2 11.0 9.8 30.2 24.1 18.8
LoRAr=32 0.83 60.5 11.7 36.1 31.7 28.0 35.4 30.4 31.4
VeRAr=16384 0.05 58.5 $\downarrow 2.0$ 9.4 $\downarrow 2.3$ 34.0 $\downarrow 2.1$ 28.9 $\downarrow 2.8$ 24.6 $\downarrow 3.4$ 35.1 $\downarrow 0.3$ 29.6 $\downarrow 0.8$ 29.5 $\downarrow 1.9$
DoRAr=32 0.84 62.3 $\uparrow 1.8$ 12.4 $\uparrow 0.7$ 37.3 $\uparrow 1.2$ 33.2 $\uparrow 1.5$ 28.8 $\uparrow 0.8$ 36.8 $\uparrow 1.4$ 31.8 $\uparrow 1.4$ 32.7 $\uparrow 1.3$
C3Ab=4096/32 0.35 61.5 $\uparrow 1.0$ 13.0 $\uparrow 1.3$ 37.2 $\uparrow 1.1$ 33.8 $\uparrow 2.1$ 29.0 $\uparrow 1.0$ 36.5 $\uparrow 1.1$ 31.9 $\uparrow 1.5$ 32.8 $\uparrow 1.4$
LLaMA3-8B Zero-Shot - 33.1 5.3 19.2 33.5 29.3 61.4 51.6 44.0
LoRAr=32 0.70 77.2 28.2 52.7 57.9 52.4 64.8 55.3 57.6
VeRAr=16384 0.04 76.0 $\downarrow 1.2$ 26.7 $\downarrow 1.5$ 51.4 $\downarrow 1.3$ 56.3 $\downarrow 1.6$ 50.8 $\downarrow 1.6$ 64.0 $\downarrow 0.8$ 53.6 $\downarrow 1.7$ 56.2 $\downarrow 1.4$
DoRAr=32 0.71 77.2 $\uparrow 0.0$ 30.0 $\uparrow 1.8$ 53.6 $\uparrow 0.9$ 59.1 $\uparrow 1.2$ 52.9 $\uparrow 0.5$ 65.7 $\uparrow 0.9$ 55.8 $\uparrow 0.5$ 58.4 $\uparrow 0.8$
C3Ab=4096/32 0.26 78.4 $\uparrow 1.2$ 29.9 $\uparrow 1.7$ 54.1 $\uparrow 1.4$ 59.5 $\uparrow 1.6$ 53.8 $\uparrow 1.4$ 65.4 $\uparrow 0.6$ 56.1 $\uparrow 0.8$ 58.7 $\uparrow 1.1$
Table 4: Comparison of various methods on the LLaMA2-7B and LLaMA3-8B models across math reasoning and code generation datasets. For each language model, the best result on each dataset is shown in bold. The symbols $\uparrow$ and $\downarrow$ indicate relative improvement and decrease, respectively, compared to the LoRA method. Experiments are conducted on a single H800 GPU with 80GB HBM.
4.1 GLUE Benchmark
Settings.
We assess our C3A on the GLUE benchmark (Wang et al., 2018) covering various tasks like single-sentence classification, similarity, paraphrase, and inference. See Table A3 in Appendix C for more details. Datasets are split into train, validation, and test sets. Models are chosen based on validation performance and evaluated on the test set. We fine-tune RoBERTa-Base and RoBERTa-Large models (Liu et al., 2019). Unique hyperparameters are from original papers (e.g., for VeRA’s $r$ and BOFT’s $b$ and $m$). Count of trainable parameters (# Params) exclude the classification head, which is uniform across methods. Shared hyperparameters (i.e., learning rates for the classification head and other parameters) are determined by hyperparameter search. Regarding the memory test, input data is fixed to 256 tokens with a batch size of 64 for consistency.
Results.